OpenStudy (anonymous):

A basketball is projected at an angle of 30 degrees. If we want the ball to go into a basket 5 meters away and +2.5 meters upwards vertically, at what speed should we shoot the ball?

OpenStudy (anonymous):

in the x or y V0x or V0y or V V means velocity

OpenStudy (anonymous):

Its asking for the velocity with the 30 degree angle... so like like vx + vy, polar form?

OpenStudy (agent0smith):

 Use \[\Large \Delta y = v_y t + 0.5 a t^2\]where delta y is your change in height, vy is your vsin30. \[\Large \Delta x = v_x t\] where vx is your vcos30. You will have to solve simultaneous equations (the t in each equation is equal)

OpenStudy (agent0smith):

Delta x is your horizontal displacement.

OpenStudy (anonymous):

Can someone actually right it all out and answer this...

OpenStudy (agent0smith):

You can at least make an attempt from what i gave. First equation is (-9.8 since gravity acts down making it negative)\[\Large 2.5 = vt \sin 30 + 0.5 (-9.8) t^2\] second equation\[\Large 5 = v t \cos 30 \]

OpenStudy (anonymous):

Thanks... I've attempted for 3 days already, still don't know how to get rid of t

OpenStudy (anonymous):

Unless I substitute one of them into the other

OpenStudy (agent0smith):

Yes, do that. Use the second equation - solve for t. Plug into the first.

OpenStudy (anonymous):

Ok, thanks... i'll try that out later... for the future, if i need to find the velocity (initial) from a projectile question i just substitute right?

OpenStudy (agent0smith):

For this type of problem, yes.